### v3

@@ 3 Conclusion
We have developed a single-shot polarization sensor based on light propagation through a short rigid piece of a few-mode fiber, resulting in a polarization-dependent scattering. This process provides a coarse speckle pattern with low signal losses and unprecedented long-term stability, which is virtually unaffected by fiber bending. The defining aspect of our approach is the sparse sampling of this pattern using a fiber array instead of capturing the whole intensity profile. This characteristic allows the use of any photonic detectors, including high-speed or single-photon detectors, enabling fast response and operation at low light intensities down to the picowatt level. The implementation of deep learning algorithms in the system ensures an accurate reconstruction, providing complete information about the polarization state, including partial polarization, with an infidelity of . Complemented by the high operation speed, it allows for both real-time and high-repetition-rate polarization sensing with remarkable precision. The all-fiber sensor, featuring a compact design with no moving components, is particularly well-suited for reducing the invasiveness of biomedical procedures in constrained spaces and under low-light conditions. Furthermore, we can enhance the spatial resolution of our sensor by tapering the tip of the few-mode fiber. This method could even provide comparable resolution as near-field scanning optical microscopy while preserving the single-photon polarization sensitivity.
@@
In summary, our polarization sensor significantly advances the field, answering the demand for a combination of compactness, accuracy, low light sensitivity, and high operational speed, which opens up new possibilities for highly accurate polarization sensing across various applications. Furthermore, our contribution lays down novel research pathways to single-photon level sensing. It also poses new questions on information and confidence evaluation of systems consisting of random media together with deep neural networks. The ability to extract meaningful data from minimal input has far-reaching implications in fields where resource efficiency and sensitivity are paramount.
4 Materials and methods
4.1 Experimental setup
@@
In addition to the information provided in the main text, we include a detailed scheme of our all-fiber polarization sensor, illustrated in Fig. 5. This scheme comprises three configurations. In the first configuration, the twisted nematic liquid crystal device [45] induces polarization transformation on the attenuated coherent signal generated by a fiber-coupled continuous 810 nm semiconductor laser diode (QPhotonics QFLD). We independently verified the accurate operation of the liquid crystal device in the second configuration by replacing it with a reference polarimeter based on rotating wave plates. The third configuration represents the sensor application and comprises the birefringent specimen under investigation. In all cases, the modified signal is collected by the 5 cm long SMF28 few-mode fiber. The number of modes supported by this fiber is characterized in Sec. 4.2.
@@
The generated speckle pattern propagates through a 5 mm free space before reaching the array of seven gradient multimode fibers GIF625. This propagation allows the speckle grain size to match the multimode fiber core. Arranged in a honeycomb structure with a 375 m diameter, the multimode fibers have a core diameter of 62.5 m and a cladding diameter of 125 m. Each fiber sparsely samples the intensity from the speckle pattern and forwards the signal to an independent single-photon avalanche diode (Excelitas SPCM). The entire experimental setup is located in an unsealed open space of a temperature-maintained laboratory.
4.2 Supported fiber modes analysis
@@
4.3 Numerical setup
The physical setup described above is assisted by a numerical processing comprising a deep neural network model. This trained network, consisting of 250 neurons per 4 hidden layers, performs a nonlinear transformation on the count distribution and returns four real-valued outputs. These represent elements of a two-by-two triangular matrix with real-valued diagonal entries and a single complex off-diagonal entry. Utilizing the Cholesky decomposition, we can reconstruct a Hermitian positive-definitive matrix Upon normalizing its trace to unity, represents a physically sound coherence matrix of the polarization state [44], mathematically equivalent to a density matrix of a two-level quantum system. Therefore, the complete model can be interpreted as providing the polarization coherence matrix of the collected light given the measured count distribution.
@@ 4.3 Numerical setup
4.4 1951 USAF birefringent test
We performed an additional analysis beyond those presented in the main text. Specifically, we conducted a scan of a 1951 USAF birefringent resolution test with the developed polarization sensor and compared it with a polarization measurement based on rotating wave plates. Fig. 6 shows a polarization image obtained with the standard method (a) alongside the scan using our all-fiber sensor (b). Both panels visualize false-colored RGB images using Bloch parameters with and corresponding to red, green, and blue channel intensities. By averaging over the background and polarization-modulating regions, we estimated a fidelity value of approximately 0.987 between the two polarization measurement approaches. This comparison unequivocally demonstrated the ability of our sensor to perform highly accurate polarization measurements.
Additionally, the image shown in panel (b) allows for the evaluation of the spatial resolution of our sensor. To this end, we fitted horizontally and vertically oriented edges at several positions. Characterized as the 20% to 80% width, we estimated the resolution to be approximately m. Besides the distance from a specimen to the fiber tip, the resolution is dictated by the few-mode fiber diameter and can be modified by employing different fiber types. Moreover, tapering the fiber tip could improve the resolution of our sensor up to the near-field scanning optical microscopy level.
Declarations
@@
The code and data that support the findings of this study are publicly available on GitHub: https://github.com/VasinkaD/Polarization-Deep-Sense.
@@
This work was supported by the Czech Science Foundation (project 21-18545S), Ministry of Education, Youth, and Sports of the Czech Republic (project OP JAC CZ.02.01.01/00/23_021/0008790). MB and DV acknowledge the support by Palacký University (project IGA-PrF-2024-008).
@@
Acknowledgements

### v4

@@ III Discussion
We have developed a single-shot polarization sensor based on light propagation through a short rigid piece of a few-mode fiber, resulting in polarization-dependent scattering. This process provides a coarse speckle pattern with low signal losses and unprecedented long-term stability, which is virtually unaffected by fiber bending. The defining aspect of our approach is the sparse sampling of this pattern using a fiber array instead of capturing the whole intensity profile. This characteristic allows the use of any photonic detectors, including high-speed or single-photon detectors, enabling fast response and operation at low light intensities down to the picowatt level. The implementation of deep learning algorithms in the system ensures an accurate reconstruction, providing complete information about the polarization state, including partial polarization, with an infidelity of . This superior performance remains stable for over a month without requiring adjustments or recalibration. Complemented by the rapid operation speed of more than two thousand high-fidelity reconstructions per second, it allows for both real-time and high-repetition-rate polarization sensing with remarkable precision. The all-fiber sensor, featuring a compact design with no moving components, is particularly well-suited for reducing the invasiveness of biomedical procedures in constrained spaces and under low-light conditions. Furthermore, one can enhance the spatial resolution of our sensor by tapering the tip of the few-mode fiber to reach the limit on near-field sensing while preserving the single-photon polarization sensitivity. However, there will be a trade-off between the increased spatial resolution and decreased collection efficiency. Also, the stability performance might be negatively affected.
@@
In summary, we presented a novel polarization sensor that is explicitly demonstrated for arbitrary polarization states, including partially polarized light, and simultaneously offers: single-photon sensitivity (picowatt powers transmitting a sample); rapid operation with thousands of reconstructions per second; high accuracy reaching Stokes errors below 0.01, and even lower at higher photon fluxes; long-term stability for over a month, unprecedented among imagers and sensors using multimode fibers; high spatial resolution of 6 m; and a compact, all-fiber design without moving components. In addition, we analyzed the resource budget of the sensor and conducted a thorough confidence analysis, which is often neglected for deep learning applications. Our polarization sensor significantly advances the field, answering the demand for a combination of compactness, accuracy, low light sensitivity, and high operational speed, which opens up new possibilities for highly accurate polarization sensing across various applications. Furthermore, our contribution lays down novel research pathways to single-photon-level sensing using random optical networks and media. It also poses new questions on information and confidence evaluation of systems consisting of random media together with deep neural networks. The ability to extract meaningful data from minimal input has far-reaching implications in fields where resource efficiency and sensitivity are paramount.
IV Materials and methods
IV.1 Experimental setup
@@
In addition to the information provided in the main text, we include a detailed scheme of our all-fiber polarization sensor, depicted in Fig. 7. This scheme comprises three configurations. In the first configuration, the twisted nematic liquid crystal device [46] induces polarization transformation on the attenuated coherent signal generated by a fiber-coupled continuous 810 nm semiconductor laser diode (QPhotonics QFLD). The laser diode exhibits typical fluctuations, drifts, and mode hops within a 1 nm wavelength range. With a more stable source, such as a distributed-feedback laser diode or solid-state laser, the wavelength stability would increase significantly, which would completely neglect any negative effects on the performance of our sensor. Moreover, the sensor can be recalibrated for an arbitrary wavelength within an hour. We independently verified the accurate operation of the liquid crystal device in the second configuration by replacing it with a reference polarimeter based on rotating wave plates. The third configuration represents the sensor application and comprises the birefringent specimen under investigation. In all cases, the modified signal is collected by the 5 cm long SMF28 few-mode fiber. The number of modes supported by this fiber is characterized in Sec. IV.2.
@@
The generated speckle pattern propagates through a 5 mm free space before reaching the array of seven gradient multimode fibers GIF625 (SQS Vláknová optika, 62.5/125/250m MM OM1 Fiber). This propagation allows the speckle grain size to match the multimode fiber core. Arranged in a honeycomb structure with a 375 m diameter, the multimode fibers have a core diameter of 62.5 m and a cladding diameter of 125 m. Each fiber sparsely samples the intensity from the speckle pattern and forwards the signal to an independent single-photon avalanche diode (Excelitas SPCM). The entire experimental setup is located in an unsealed open space of a temperature-maintained laboratory, with the temperature varying approximately within 1.5 °C and the relative humidity ranging from 25% to 35% during the measurements. Despite these fluctuations, we demonstrated unprecedented stability and accuracy over many weeks (see Section II.2). This level of robustness is fully sufficient for any laboratory measurements. Moreover, in typical biomedical and material research, the environmental conditions, such as sample temperature, are deliberately controlled and stabilized. Under such conditions, the long-term stability of our sensor will be even greater than demonstrated.
IV.2 Supported fiber modes analysis
@@
IV.3 Numerical setup
The physical setup described above is assisted by a numerical processing comprising a deep neural network model. This trained network, consisting of 250 neurons per 4 hidden layers, performs a nonlinear transformation on the count distribution and returns four real-valued outputs. These represent elements of a two-by-two triangular matrix with real-valued diagonal entries and a single complex off-diagonal entry. Utilizing the Cholesky decomposition, we can reconstruct a Hermitian positive-definitive matrix Upon normalizing its trace to unity, represents a physically sound coherence matrix of the polarization state [45], mathematically equivalent to a density matrix of a two-level quantum system. Therefore, the complete model can be interpreted as providing the polarization coherence matrix of the collected light given the measured count distribution. Alternatively, the polarization state can be described by the Bloch parameters , , where are Pauli matrices.
@@ IV.3 Numerical setup
IV.4 1951 USAF birefringent test
We conducted a scan of a 1951 USAF birefringent resolution test with the developed polarization sensor and compared it with a polarization measurement based on rotating wave plates. Fig. 8 shows a polarization image obtained with the standard method (a) alongside the scan using our all-fiber sensor (b). Both panels visualize false-colored RGB images using Bloch parameters with and corresponding to red, green, and blue channel intensities. By averaging over the background and polarization-modulating regions, we estimated a fidelity value of approximately 0.987 between the two polarization measurement approaches. This comparison unequivocally demonstrated the ability of our sensor to perform highly accurate polarization measurements.
Additionally, the image shown in panel (b) allows for the evaluation of the spatial resolution of our sensor. To this end, we fitted horizontally and vertically oriented edges at several positions. Characterized as the 20% to 80% width, we estimated the resolution to be approximately m. Besides the distance from a specimen to the fiber tip, the resolution is dictated by the few-mode fiber diameter and can be modified by employing different fiber types. Moreover, for applications requiring higher resolution, utilizing fiber tip tapering could improve it up to the near-field scanning optical microscopy at the cost of the collection efficiency and possibly stability.
Acknowledgements
@@
The code and data that support the findings of this study are publicly available on GitHub:
https://github.com/VasinkaD/Polarization-Deep-Sense.
@@
This work was supported by the Czech Science Foundation (project 21-18545S), Ministry of Education, Youth, and Sports of the Czech Republic (project OP JAC CZ.02.01.01/00/23_021/0008790). MB and DV acknowledge the support by Palacký University (projects IGA-PrF-2024-008 and IGA-PrF-2025-010).
@@
Conflict of Interest
The authors declare that they have no conflict of interest.